# Beyond Local Sharpness: Communication-Efficient Global Sharpness-aware Minimization

### v1

@@ 3.3 SAM in Federated Learning
FedSam (Caldarola et al., 2022; Qu et al., 2022) aims to improve the clients’ models generalization through convergence to flatter regions by using Sam in the local training. From Eqs. 1 and 4, the global objective becomes $\min_{\operatorname{\boldsymbol{w}}}\left\{f^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})=\nicefrac{{1}}{{C}}\sum_{k\in\operatorname{\mathcal{C}}}f_{k}^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})\right\}$, with $f_{k}^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})\triangleq\max_{\|\operatorname{\boldsymbol{\epsilon}}_{k}\|\leq\rho}f_{k}(\operatorname{\boldsymbol{w}}+\operatorname{\boldsymbol{\epsilon}}_{k})$ with local perturbation $\operatorname{\boldsymbol{\epsilon}}_{k}$. The intuition behind this approach is that the improved local models’ generalization positively reflects on the global model performance. However, by independently applying Eq. 4 in the local optimization, FedSam does not explicitly address global flatness, potentially leading to discrepancies between local and global loss geometries.
@@ 4 Local-Global Sharpness Inconsistency
This section empirically investigates the hypothesis that discrepancies between local and global loss landscapes impact FedSam’s performance, using a CNN model on Cifar10 and Cifar100 datasets — further details in Section 6.
Fig. 1 compares the loss surfaces of CNNs trained with FedAvg and FedSam. On the easier Cifar10, FedSam exhibits noticeably flatter minima w.r.t. FedAvg, effectively navigating simpler landscapes. However, their performance difference diminishes with increasing dataset complexity (Cifar100) and heterogeneity ($\alpha\rightarrow 0$). This suggests that larger discrepancies between local and global geometries arise as tasks become more complex and data distributions more diverse.
@@
To highlight the existing difference between local and global behavior, Fig. 2 investigates the behavior of client models at the end of local training when tested on their own data $\operatorname{\mathcal{D}}_{k}$ (bottom landscape), prior to server-side aggregation, w.r.t. the overall dataset $\operatorname{\mathcal{D}}$ (top landscape). Each plot shows the behavior of one of the randomly selected clients during the last round with FedSam, distinguished by the locally seen class (results for all 5 clients in Appendix B). The inconsistency between local and global behavior can be easily appreciated: locally, each model lands in a flat region; differently, the same model is close to saddle points (Fig. 2(a)) or sharp minima (Figs. 2(b) and 2(c)) in the global landscape. These findings are further corroborated by the Hessian eigenvalues presented in Table 1. FedSam’s local maximum Hessian eigenvalue, denoted by $\lambda_{1,l}$ and computed on each client’s individual dataset, is significantly lower than the global eigenvalue $\lambda_{1,g}$, calculated on the overall dataset, on the more complex Cifar100. This suggests that FedSam effectively achieves local convergence to flatter regions of the loss landscape on individual devices. However, the higher global eigenvalue indicates limitations in reaching a globally flat minimum. The challenge of achieving flat regions under high heterogeneity and the gap between local and global flatness support the introduction of FedGloSS.
Table 1: Maximum Hessian eigenvalues of local models, computed on global ($\lambda_{1,g}$) and local datasets ($\lambda_{1,l}$). Cifar10 and Cifar100, $\alpha=0$. Each client is identified via its local class. The lowest $\lambda_{1,g}$ in bold. FedDyn does not converge on Cifar100 with $\alpha=0$ (Caldarola et al., 2022; Varno et al., 2022), hence the lack of results ().
@@ 5.1 Rethinking SAM in Federated Learning
Aiming to optimize Sam’s objective (Eq. 4) on the global function, FedGloSS solves $\min_{\operatorname{\boldsymbol{w}}}\left\{\operatorname{\mathcal{F}}(\operatorname{\boldsymbol{w}})=\frac{1}{C}\sum_{k\in\operatorname{\mathcal{C}}}\operatorname{\mathcal{F}}_{k}(\operatorname{\boldsymbol{w}})\right\}$, with $\operatorname{\mathcal{F}}_{k}(\operatorname{\boldsymbol{w}})\triangleq\max_{\|\operatorname{\boldsymbol{\epsilon}}\|\leq\rho}f_{k}(\operatorname{\boldsymbol{w}}+\operatorname{\boldsymbol{\epsilon}})$, where $\operatorname{\boldsymbol{\epsilon}}$ is the global perturbation. Calculating the true $\operatorname{\boldsymbol{\epsilon}}$ value requires the global gradient $\nabla_{\operatorname{\boldsymbol{w}}}f$ (Eq. 5) computed on the entire dataset $\operatorname{\mathcal{D}}\triangleq\cup_{k\in\operatorname{\mathcal{C}}}\operatorname{\mathcal{D}}_{k}$, which is not available in FL due to data privacy and communication constraints. While FedSmoo (Sun et al., 2023a) tackles this issue by using ADMM on the sharpness with the constraint $\operatorname{\boldsymbol{\epsilon}}=\operatorname{\boldsymbol{\epsilon}}_{k}$, it necessitates transmitting $\operatorname{\boldsymbol{\epsilon}}$ alongside the model parameters $\operatorname{\boldsymbol{w}}$ to both clients and server in each round, hindering its practicality in real-world scenarios with limited communication budgets. This observation motivates the question: how to minimize global sharpness while maintaining communication efficiency?
@@ 5.2 FedGloSS
FedSam (Caldarola et al., 2022; Qu et al., 2022) Sgd Sam $1\times$ $2\times$
FedDyn (Acar et al., 2021) + Sam Sgd Sam $1\times$ $2\times$
FedSpeed (Sun et al., 2023b) Sgd Similar to Sam $1\times$ $2\times$
FedGamma (Dai et al., 2023) Sgd Sam $2\times$ $2\times$
FedSmoo (Sun et al., 2023a) Sgd Sam $2\times$ $2\times$
@@
To overcome the challenges posed by NaiveFedGloSS, following (Park et al., 2023), FedGloSS estimates $\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}$ using the perturbed global pseudo-gradient from the previous round $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ at each round $t$. This approach leverages available information without incurring extra communications and avoiding unnecessary computations. Intuitively, the use of the previous pseudo-gradient to minimize the sharpness allows FedGloSS to access information on the global loss landscape geometry, thus guiding the global optimization towards flatter minima. From Eqs. 5 and 3, FedGloSS updates the global model $\operatorname{\boldsymbol{w}}^{t}$ as
@@
[…]
where with a slight abuse of notation ServerOpt from Eq. 3 is substituted with the server-side strategy proposed by FedGloSS. The notation follows the colors of Fig. 3, which depicts our approach. Notably, as summarized in Table 2, FedGloSS enables Sam on the server side while allowing any ClientOpt for local training, with computational costs varying based on the chosen optimizer. This differs from previous methods constrained to the more computationally expensive Sam. In addition, differently from FedSmoo, FedGloSS maintains FedAvg’s communication complexity while optimizing for global flatness.
@@ 5.2.1 Promoting global consistency with ADMM
$\delta_{\epsilon}^{t}\triangleq\|\operatorname{\tilde{\boldsymbol{\epsilon}}}^{t}(\operatorname{\boldsymbol{w}})-\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}(\operatorname{\boldsymbol{w}})\|=\rho\left\|\frac{\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}}{\|\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}\|}-\frac{{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}}{\|{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}\|}\right\|,$ (6)
where $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ is computed using the updates of the clients in $\operatorname{\mathcal{C}}^{t-1}$ and $\tilde{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}$ with $\operatorname{\mathcal{C}}^{t}$. Eq. 6 suggests $\delta_{\epsilon}^{t}$ is minimized when $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ and $\tilde{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}$ are aligned. However, in real-world heterogeneous FL, i) to due clients’ unavailability, only a subset of them participates in training at each round, with $\operatorname{\mathcal{C}}^{t}$ likely differing from $\operatorname{\mathcal{C}}^{t-1}$, and ii) clients hold different data distributions, i.e., local optimization paths likely converge towards different local minima, leading to unstable global updates (Karimireddy et al., 2020b). As a consequence, $\delta_{\epsilon}^{t}\not\to 0$ necessarily.
To align local and global objectives - ensuring client and server gradient alignment and minimizing Eq. 6 - FedGloSS leverages the Alternating Direction Method of Multipliers (ADMM) (Boyd et al., 2011) on $\operatorname{\boldsymbol{w}}^{t}$ (Acar et al., 2021; Sun et al., 2023a; b). While alternative approaches could be used, they either lack full immunity to data heterogeneity or have shown poor performance on realistic scenarios (e.g., variance reduction Karimireddy et al. (2020b); Dai et al. (2023)). In contrast, ADMM has been proved to converge under arbitrary heterogeneity Acar et al. (2021) and can thus be leveraged as a base algorithm for FedGloSS, as shown in Algorithm 1 in Appendix A. ADMM makes use of the augmented Lagrangian function $\mathcal{L}(\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{W}},\sigma)=\sum_{k\in\operatorname{\mathcal{C}}}L(\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{w}}_{k},\sigma_{k})$ where $\operatorname{\boldsymbol{W}}=\{\operatorname{\boldsymbol{w}}_{1},\cdots,\operatorname{\boldsymbol{w}}_{C}\}$ and $\sigma$ is the Lagrangian multiplier. The problem solved by $\mathcal{L}$ is
@@
with $\beta>0$ being an hyperparameter. Eq. 7 is split into $C$ sub-problems of the form $\operatorname{\boldsymbol{w}}_{k,E}=\operatorname*{arg\,min}_{\operatorname{\boldsymbol{w}}_{k}}\{f_{k}-\sigma_{k}^{\top}(\operatorname{\boldsymbol{w}}^{t}-\operatorname{\boldsymbol{w}}_{k})+\frac{1}{2\beta}\|\operatorname{\boldsymbol{w}}^{t}-\operatorname{\boldsymbol{w}}_{k}^{t}\|^{2}\}$. The local dual variable is updated as $\sigma_{k}\leftarrow\sigma_{k}-\frac{1}{\beta}(\operatorname{\boldsymbol{w}}_{k,E}^{t}-\operatorname{\boldsymbol{w}}_{k,0}^{t})$. The global one $\sigma$ is updated by adding the averaged $\operatorname{\boldsymbol{w}}_{k}-\operatorname{\boldsymbol{w}}^{t}\,\forall k\in\operatorname{\mathcal{C}}$.
Figure 4 confirms the effect of ADMM on gradient alignment: the difference between the true and approximated perturbation, $\delta_{\epsilon}^{t}$ (Eq. 6), decreases over training rounds and with the use of Lagrangian multipliers.
@@ 6.1 Experimental setting
Appendix C details implementation and hyperparameter settings.
Federated datasets. We leverage established FL benchmarks (Caldas et al., 2019; Hsu et al., 2020; 2019). Small-scale image classification: following (Hsu et al., 2020; Caldarola et al., 2022), the federated versions of Cifar10 (10 classes) and Cifar100 (100 classes) (Krizhevsky, 2009) split the respective $50k$ training images in 100 clients with 500 images each. The data distribution is controlled by the Dirichlet’s parameter $\alpha\in\{0,0.05\}$ for Cifar10 and $\{0,0.5\}$ for Cifar100 (Hsu et al., 2019). Lower $\alpha$ signifies increased heterogeneity, with $\alpha=0$ being the most challenging scenario where each client holds samples from one class. Large-scale image classification: Landmarks-User-160k ($2,028$ classes) (Hsu et al., 2020) is the federated Google Landmarks v2 (Weyand et al., 2020) with $164,172$ pictures of worldwide locations, split among $1,262$ realistic clients.
Models. The effectiveness of FedGloSS is shown using multiple model architectures. As in (Hsu et al., 2020; Caldarola et al., 2022), we use a Convolutional Neural Network (CNN) similar to LeNet5 (LeCun et al., 1998) on both Cifar10 ($T=10k$) and Cifar100 ($T=20k$). Experiments with ResNet18 (He et al., 2015) run for $10k$ rounds. For Landmarks-User-160k, we train MobileNetv2 (Sandler et al., 2018; Hsu et al., 2020) ($T=1.3k$), considering the limited resources at the edge.
Baselines. To study real-world settings with varying participation, a small fraction of clients is sampled at each round, with participation rate set to $5\%$ with the CNN and $10\%$ with ResNet18 on both Cifars, and to 50 clients per round in Landmarks-User-160k ($\approx 4\%$). FedGloSS is compatible with any local optimizer (Section 5). We choose SGD and SAM to comply with previous works and compare it with state-of-the-art (SOTA) methods for statistical heterogeneity in FL, distinguishing the results by optimizer type to highlight performance differences. Sgd-based approaches are FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020c), FedDyn (Acar et al., 2021) and Scaffold (Karimireddy et al., 2020b), while use Sam FedSAM (Caldarola et al., 2022; Qu et al., 2022), FedDyn + SAM, FedSpeed (Sun et al., 2023b), FedGamma (Dai et al., 2023) and FedSmoo (Sun et al., 2023a).

### v2

@@ 3.3 SAM in Federated Learning
FedSam [4, 45] aims to improve the clients’ models generalization through convergence to flatter regions by using Sam in the local training. From Eqs. 1 and 4, the global objective becomes $\min_{\operatorname{\boldsymbol{w}}}\left\{f^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})=\nicefrac{{1}}{{C}}\sum_{k\in\operatorname{\mathcal{C}}}f_{k}^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})\right\}$, with $f_{k}^{\textsc{Sam}}(\operatorname{\boldsymbol{w}})\triangleq\max_{\|\operatorname{\boldsymbol{\epsilon}}_{k}\|\leq\rho}f_{k}(\operatorname{\boldsymbol{w}}+\operatorname{\boldsymbol{\epsilon}}_{k})$ with local perturbation $\operatorname{\boldsymbol{\epsilon}}_{k}$. The intuition behind this approach is that the improved local models’ generalization positively reflects on the global model performance. However, by independently applying Eq. 4 in the local optimization, FedSam does not explicitly address global flatness, potentially leading to discrepancies between local and global loss geometries.
@@ 4 Local-Global Sharpness Inconsistency
This section empirically investigates the hypothesis that discrepancies between local and global loss landscapes impact FedSam’s performance, using a CNN model on Cifar10 and Cifar100 datasets — further details in Sec. 6.
Fig. 1 compares the loss surfaces of CNNs trained with FedAvg and FedSam. On the easier Cifar10, FedSam exhibits flatter minima w.r.t. FedAvg, effectively navigating simpler landscapes. However, their difference diminishes with increasing dataset complexity (Cifar100) and heterogeneity ($\alpha\rightarrow 0$). This suggests larger discrepancies between local and global geometries arise as tasks become more complex and data distributions more diverse.
To highlight the existing difference between local and global behavior, Fig. 2 investigates the behavior of client models at the end of local training when tested on their own data $\operatorname{\mathcal{D}}_{k}$ (bottom landscape), prior to server-side aggregation, w.r.t. the overall dataset $\operatorname{\mathcal{D}}$ (top landscape). Each plot shows the behavior of one of the randomly selected clients during the last round with FedSam, distinguished by the locally seen class (results for all 5 clients in App. B). The inconsistency between local and global behavior can be easily appreciated: locally, each model lands in a flat region; differently, the same model is close to saddle points (Fig. 2(a)) or sharp minima (Figs. 2(b) and 2(c)) in the global landscape. These findings are further corroborated by the maximum Hessian eigenvalues presented in Tab. 1, computed using each client’s local dataset ($\lambda_{1,l}$) and the overall one ($\lambda_{1,g}$). FedSam performs well on simple datasets like Cifar10 with global eigenvalues smaller than local ones. However, $\lambda_{1,l}\ll\lambda_{1,g}$ on the more complex Cifar100. This suggests that FedSam effectively achieves local convergence to flatter regions of the loss landscape on individual devices, but the higher global eigenvalue indicates limitations in reaching a globally flat minimum. The challenge of achieving flat regions under high heterogeneity and the gap between local and global flatness support the introduction of FedGloSS.
@@
Table 1: Maximum Hessian eigenvalues of local models, computed on global ($\lambda_{1,g}$) and local datasets ($\lambda_{1,l}$). Cifar10 and Cifar100, $\alpha=0$. Each client is identified via its local class. The lowest $\lambda_{1,g}$ in bold. FedDyn does not converge on Cifar100 with $\alpha=0$ [4, 56], hence the lack of results ().
@@ 5.1 Rethinking SAM in Federated Learning
Aiming to optimize Sam’s objective (Eq. 4) on the global function, FedGloSS solves $\min_{\operatorname{\boldsymbol{w}}}\left\{\operatorname{\mathcal{F}}(\operatorname{\boldsymbol{w}})=\frac{1}{C}\sum_{k\in\operatorname{\mathcal{C}}}\operatorname{\mathcal{F}}_{k}(\operatorname{\boldsymbol{w}})\right\}$, with $\operatorname{\mathcal{F}}_{k}(\operatorname{\boldsymbol{w}})\triangleq\max_{\|\operatorname{\boldsymbol{\epsilon}}\|\leq\rho}f_{k}(\operatorname{\boldsymbol{w}}+\operatorname{\boldsymbol{\epsilon}})$, where $\operatorname{\boldsymbol{\epsilon}}$ is the global perturbation. Calculating the true $\operatorname{\boldsymbol{\epsilon}}$ value requires the global gradient $\nabla_{\operatorname{\boldsymbol{w}}}f$ (Eq. 5) computed on the entire dataset $\operatorname{\mathcal{D}}\triangleq\cup_{k\in\operatorname{\mathcal{C}}}\operatorname{\mathcal{D}}_{k}$, which is not available in FL due to data privacy and communication constraints. While FedSmoo [53] tackles this issue by using ADMM on the sharpness with the constraint $\operatorname{\boldsymbol{\epsilon}}=\operatorname{\boldsymbol{\epsilon}}_{k}$, it necessitates transmitting $\operatorname{\boldsymbol{\epsilon}}$ alongside the model parameters $\operatorname{\boldsymbol{w}}$ to both clients and server in each round, hindering its practicality in real-world scenarios with limited communication budgets. This observation motivates the question: how to minimize global sharpness while maintaining communication efficiency?
@@ 5.1.1 Challenges of Server-side SAM
FedSam [4, 45] Sgd Sam $1\times$ $2\times$
FedDyn [1] + Sam Sgd Sam $1\times$ $2\times$
FedSpeed [54] Sgd Similar to Sam $1\times$ $2\times$
FedGamma [7] Sgd Sam $2\times$ $2\times$
FedSmoo [53] Sgd Sam $2\times$ $2\times$
@@
5.2 FedGloSS
To overcome the challenges posed by NaiveFedGloSS, following [42], FedGloSS estimates $\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}$ using the perturbed global pseudo-gradient from the previous round $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ at each round $t$. This approach leverages available information without incurring extra communications and avoids unnecessary computations. Intuitively, the use of the previous pseudo-gradient to minimize the sharpness allows FedGloSS to access information on the global loss landscape geometry, thus guiding the global optimization towards flatter minima. From Eqs. 5 and 3, FedGloSS updates the global model $\operatorname{\boldsymbol{w}}^{t}$ as
@@ 5.2 FedGloSS
where with a slight abuse of notation ServerOpt from Eq. 3 is substituted with the server-side strategy proposed by FedGloSS. The notation follows the colors of Fig. 3, which depicts our approach. Notably, as summarized in Tab. 2, FedGloSS enables Sam on the server side while allowing any ClientOpt for local training, with computational costs varying based on the chosen optimizer. This differs from previous methods constrained to the more computationally expensive Sam. In addition, differently from FedSmoo, FedGloSS maintains FedAvg’s communication complexity while optimizing for global flatness.
(a) Local model on class sea
w/ FedGloSS
(b) Local model on class snail
w/ FedGloSS
(c) Local model on class sea
w/ FedSmoo
(d) Local model on class snail
w/ FedSmoo
@@ 5.2.1 Promoting Global Consistency with ADMM
$\vspace{-5pt}\delta_{\epsilon}^{t}\triangleq\|\operatorname{\tilde{\boldsymbol{\epsilon}}}^{t}(\operatorname{\boldsymbol{w}})-\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}(\operatorname{\boldsymbol{w}})\|=\rho\left\|\frac{\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}}{\|\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}\|}-\frac{{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}}{\|{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}\|}\right\|,$ (6)
where $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ is computed using the updates of the clients in $\operatorname{\mathcal{C}}^{t-1}$ and $\tilde{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}$ with $\operatorname{\mathcal{C}}^{t}$. Eq. 6 suggests $\delta_{\epsilon}^{t}$ is minimized when $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ and $\tilde{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}$ are aligned, which occurs when clients’ updates are directionally consistent. However, in real-world heterogeneous FL, i) to due clients’ unavailability, only a subset of them participates in training at each round, with $\operatorname{\mathcal{C}}^{t}$ likely differing from $\operatorname{\mathcal{C}}^{t-1}$, and ii) clients hold different data distributions, i.e., local optimization paths likely converge towards different local minima, leading to unstable global updates [26]. As a consequence, $\delta_{\epsilon}^{t}\not\to 0$ necessarily.
To align local and global objectives - guiding client and server updates in the same direction and minimizing Eq. 6 - FedGloSS leverages the Alternating Direction Method of Multipliers (ADMM) [3] on $\operatorname{\boldsymbol{w}}^{t}$ [1, 53, 54]. While alternative approaches could be used, they either lack full immunity to data heterogeneity or have shown poor performance on realistic scenarios (e.g., variance reduction [26, 7]). In contrast, ADMM has been proved to converge under arbitrary heterogeneity [1] and can thus be leveraged as a base algorithm for FedGloSS, as shown in Alg. 1 in App. A. ADMM makes use of the augmented Lagrangian function $\mathcal{L}(\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{W}},\sigma)=\sum_{k\in\operatorname{\mathcal{C}}}L(\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{w}}_{k},\sigma_{k})$ where $\operatorname{\boldsymbol{W}}=\{\operatorname{\boldsymbol{w}}_{1},\cdots,\operatorname{\boldsymbol{w}}_{C}\}$ and $\sigma$ is the Lagrangian multiplier. The problem solved by $\mathcal{L}$ is
@@
with $\beta>0$ being an hyperparameter. Eq. 7 is split into $C$ sub-problems of the form $\operatorname{\boldsymbol{w}}_{k,E}=\operatorname*{argmin}_{\operatorname{\boldsymbol{w}}_{k}}\{f_{k}-\sigma_{k}^{\top}(\operatorname{\boldsymbol{w}}^{t}-\operatorname{\boldsymbol{w}}_{k})+\frac{1}{2\beta}\|\operatorname{\boldsymbol{w}}^{t}-\operatorname{\boldsymbol{w}}_{k}^{t}\|^{2}\}$. The local dual variable is updated as $\sigma_{k}\leftarrow\sigma_{k}-\frac{1}{\beta}(\operatorname{\boldsymbol{w}}_{k,E}^{t}-\operatorname{\boldsymbol{w}}_{k,0}^{t})$. The global one $\sigma$ is updated by adding the averaged $\operatorname{\boldsymbol{w}}_{k}-\operatorname{\boldsymbol{w}}^{t}\,\forall k\in\operatorname{\mathcal{C}}$. We note that ADMM introduces a key limitation of our approach, namely reliance on stateful clients. Fig. 5 confirms the directional consistency of local and global updates enforced by ADMM reduces $\delta_{\epsilon}^{t}$ (Eq. 6), i.e., the difference between the true and approximated perturbation. The gap between the directions of $\tilde{\Delta}^{t-1}_{\operatorname{\boldsymbol{w}}}$ and $\tilde{\Delta}^{t}_{\operatorname{\boldsymbol{w}}}$ remains constant after an initial phase, suggesting the most challenging loss landscape direction is largely stable over time.
@@ 6.1 Experimental Setting
App. C details implementation and hyperparameter settings.
Federated datasets. We leverage established FL benchmarks [6, 20, 19]. Small-scale image classification: following [20, 4], the federated versions of Cifar10 (10 classes) and Cifar100 (100 classes) [29] split the respective $50k$ training images in 100 clients with 500 images each. The data distribution is controlled by Dirichlet’s parameter $\alpha\in\{0,0.05,1,5,10\}$ for Cifar10 and $\{0,0.5\}$ for Cifar100 [19]. Lower $\alpha$ signifies increased heterogeneity, with $\alpha=0$ being the most challenging scenario (each client holds samples from one class). Large-scale image classification: Landmarks-User-160k ($2,028$ classes) [20] is the federated Google Landmarks v2 [58] with $164,172$ pictures of worldwide locations, split among $1,262$ realistic clients.
Models. The effectiveness of FedGloSS is shown using multiple model architectures. As in [20, 4], we use a Convolutional Neural Network (CNN) similar to LeNet5 [30] on both Cifar10 ($T=10k$) and Cifar100 ($T=20k$). Experiments with ResNet18 [16] run for $10k$ rounds. For Landmarks-User-160k, we train MobileNetv2 [50, 20] ($T=1.3k$), considering the limited resources at the edge.
Baselines. To study real-world settings with varying participation, a small fraction of clients is sampled at each round, with participation rate set to $5\%$ with the CNN and $10\%$ with ResNet18 on both Cifars, and to 50 clients per round in Landmarks-User-160k ($\approx 4\%$). FedGloSS is compatible with any local optimizer (Sec. 5). We choose SGD and SAM to comply with previous works and compare it with state-of-the-art (SOTA) methods for statistical heterogeneity in FL, distinguishing the results by optimizer type to highlight performance differences. Sgd-based approaches are FedAvg [39], FedProx [35], FedDyn [1] and Scaffold [26], while use Sam FedSAM [4, 45], FedDyn + SAM, FedSpeed [54], FedGamma [7] and FedSmoo [53].